Money and Mixture Problems:
Jerry invested $7000, part at 8% simple interest and the rest at 5% simple interest for a period of 1 year. If he received a total annual interest of $476 from both investments, how much did he invest at each rate?

Answer by mananth(16946)

Investment I 8.00% per annum
Investment II 5.00% per annum

x+y 7000 ------------------------1
8.00%x+5.00%y=$476.00
Multiply by 100
8x+5y= $47,600.00 --------2

Multiply (1) by -8
we get

-8x-8y= -56000.00

Add this to (2)

-3y=-$8,400.00

divide by-3

y=$2,800.00----investment at 5.00%
Balance $4,200.00----investment at 8.00%
m.ananth@hotmail.ca